OpenStudy (anonymous):

You decide to put $100 in a savings account to save for a $3,000 down payment on a new car. If the account has an interest rate of 2% per year and is compounded monthly, how long does it take you to earn $3,000 without depositing any additional funds?

OpenStudy (anonymous):

@ospreytriple

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

Do you know the compound interest formula from your course?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

a(t)=p(1+r)^t

OpenStudy (anonymous):

OK. It looks like we want a(t) to be $3000 cause that's how much needs to be saved. p is the present value. How much is that?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

OH!! In your previous question, the right hand side of the equation was NEGATIVE, so the correct answer is the negative number. Sorry about that.

OpenStudy (anonymous):

$100 is correct.

OpenStudy (anonymous):

The interest rate (r) is a little bit tricky. The question says the nominal yearly interest rate is 2% but it is compounded monthly. Do you know how to get the interest rate in this case?

OpenStudy (anonymous):

ummm 12??

OpenStudy (anonymous):

Well, it's 2% per year, but for a month it's going to be 1/12 as much. What do you think?

OpenStudy (anonymous):

I agree

OpenStudy (anonymous):

OK. so r = 2% /12. Have to convert that to a decimal. What do you get?

OpenStudy (anonymous):

0.02

OpenStudy (anonymous):

Right. But you still have to divide by 12 to get r.

OpenStudy (anonymous):

ok so 0.02/12 ??

OpenStudy (mertsj):

\[A=P(1+\frac{r}{n})^{nt}\] \[3000=100(1+\frac{.02}{12})^{12t}\]

OpenStudy (mertsj):

Solve for t

OpenStudy (anonymous):

That's it. What's r then?

OpenStudy (anonymous):

3000=100.16667)^12t

OpenStudy (anonymous):

r is not 0.16667. Try it again

OpenStudy (anonymous):

r=0.02

OpenStudy (anonymous):

Remember, we already established that r = 0.02/12. What do you get?

OpenStudy (anonymous):

0.001666667

OpenStudy (anonymous):

Great. Now out that value of r into the compound interest formula and what do you get?

OpenStudy (mertsj):

r=.02

OpenStudy (anonymous):

Perhaps you'd like to take over @Mertsj

OpenStudy (anonymous):

3000=(100.16667)^12t

OpenStudy (anonymous):

What's 100 + 0.0016667 ?

OpenStudy (mertsj):

Continuing: \[\frac{3000}{100}=\frac{100(1+\frac{.02}{12})^{12t}}{100}\]

OpenStudy (anonymous):

100.0016667

OpenStudy (anonymous):

30=(1+0.02/12)^12t

OpenStudy (anonymous):

Great. So you have\[3000 = 100.0016667^{12t}\]Now take the log of both sides

OpenStudy (mertsj):

\[300=(1.01666666)^{12t}\]

OpenStudy (anonymous):

ok so 1og 300=....?

OpenStudy (mertsj):

\[\log_{10}300=12t \log_{10}1.0166666666 \]

OpenStudy (anonymous):

I'm sorry @BellaBlue77 . Our work is being hijacked. I'm out. Good luck.

OpenStudy (mertsj):

\[\frac{\log_{10}300}{\log_{10}1.016666666}=12t \]

OpenStudy (anonymous):

ok hold onn

OpenStudy (anonymous):

300/ 1.01666666=12t

OpenStudy (mertsj):

\[345.07=12t\]

OpenStudy (anonymous):

28. 76

OpenStudy (mertsj):

\[28.8=t\]

OpenStudy (anonymous):

170.202 years 14.3129 years 171.755 years 168.354 years

OpenStudy (anonymous):

That's not an answer choice

OpenStudy (anonymous):

@heretohelpalways @ganeshie8

OpenStudy (mertsj):

Error: 3000/100=30

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[\frac{\log_{10}30}{\log_{10}1.0166666666}=12t \]

OpenStudy (anonymous):

ok so 30/1.016666

OpenStudy (anonymous):

29.5081967232=12t

OpenStudy (mertsj):

you have to take the log of those numbers.

OpenStudy (anonymous):

so log 29.508...?

OpenStudy (mertsj):

Do you see that it says log 30?

OpenStudy (anonymous):

1.46994267012

OpenStudy (anonymous):

1.47712125472 i mean

OpenStudy (anonymous):

would 168.354 years be the answer?

OpenStudy (mertsj):

I keep getting 17.14

OpenStudy (anonymous):

I know :( But thats not an answer choice. 170.202 years 14.3129 years 171.755 years 168.354 years see?

OpenStudy (anonymous):

Oh well, let's call it a day loll

OpenStudy (anonymous):

If you find the answer, message me pleae, I'll be back tommorw

OpenStudy (mertsj):

Starting over: \[3000=100(1+\frac{.02}{12})^{12t}\] \[30=(1+.00166666666)^{12t}\]

OpenStudy (mertsj):

\[\frac{\log_{10}30}{\log_{10}1.00166666666666}=12t \]

OpenStudy (mertsj):

\[2042.4185=12t\]

OpenStudy (mertsj):

\[170.20=t\]